# Statistical Validation of Contagion Centrality in Financial Networks

### v1

@@ 4.2 Equity Tick Data
The choice of data frequency is a critical consideration in network analysis. To ensure that dependencies are captured in the model, we must strike the right balance between a frequency that is not too low or too high. Given that stock prices swiftly adjust to market changes, especially to similar stock price fluctuations, and trading occurs at high frequency to capitalize on these changes, we opt for tick data. We use tick log return data for 20 Financials sector companies (provided in Table 3 of Appendix) as well as both Lehman Brothers (LEH) and Bear Stearns (BSC). The tick data is obtained from Refinitiv, collected at one-second intervals, and spans from January 1, 2005 to December 31, 2010.
@@
In Figure 7, we focus on the period surrounding the 2008 Financial Crisis to gain a deeper understanding of the centrality measure’s behavior in the years preceding, during, and after the turmoil. We consider two versions of the system-level LCM: (i) for the network including LEH and BSC, for which we run the analysis for as long as the data is available (October 8, 2008 indicated by a red dashed line in the figure) and (ii) for the network excluding these two firms. In both versions, we observe an upward trend starting from early 2007 with the peak LCM reached in August 2008. A gradual decline is then observed when excluding LEH and BSC. This signifies that the contagion risk increased until reaching the peak of turmoil, subsequently decreasing after government interventions and as the economy and markets slowly returned to normalcy. This system-level trend closely aligns with the timeline of the financial crisis: the acquisition of Bear Sterns by JPMorgan on March 16, 2008, and the subsequent default of Lehman Brothers on September 15, 2008. The Wall Street Bailout law was passed on October 3, 2008, which included the \$700 billion Troubled Asset Relief Program (TARP) aimed at rescuing banks from defaulting.
@@
We plot the node-level LCM for three specific firms (American International Group (AIG), JPMorgan (JPM) and Goldman Sachs (GS)) alongside Lehman Brothers and Bear Stearns in Figure 7. We note that the node-level LCM for LEH and BSC have similar behavior to the system-level LCM until May 2008. Afterwards, they start to decline, due to which we experience a slight decline in the system-level LCM as well. The node-level drop is aligned with recognition of the substantial risk these two companies carried by investors. Due to this market behavior, investors treated these companies idiosyncratically thus reducing the spillover effects to the rest of the system from shocks they might experience. This is why, despite the systemic contagion risk rising, the LCM for LEH and BSC move inversely. This detachment is also reflected in the network structure. Figure 8 shows the interdependencies between LEH and BSC and all other firms for mid March and September 2008 time windows. As is evident, the connectivity of these two companies diminishes in both quantity and significance during their corresponding drop in their LCM.
@@
Analyzing node-level LCMs reveals some similarities and differences between the three tickers, AIG, JPM and GS, and the delisted ones. We notice that AIG, prominent in the insurance sector, exhibits a similar behavior to LEH and BSC; its LCM took a downward turn following the bankruptcy and merger of LEH and BSC respectively. During the crisis, \$182 billion was injected into AIG to prevent its default by the Treasury and Federal Reserve Bank of New York, as it was believed that its collapse would jeopardize the integrity of other firms. Essentially, AIG came under government ownership and was safeguarded, which is reflected in its LCM as it reaches nearly zero by early 2009. On the other hand, JPM and GS, for which the LCM consistently increases even after October 8, 2008, diverge from the system-level LCM. As these two companies were considered safer by investors, they maintained investments despite the systemic turmoil. However, under a shock scenarios to these companies, there would have been a high contagion risk. Looking into the stock prices of these tickers, we observe a significant decline in AIG’s price, whereas JPM and GS experience a decrease but recover shortly afterwards. These patterns are aligned with the fluctuations seen in the LCM, indicating its ability to mirror the dynamics of the market.
Figure 7: Node- and system-level LCM
@@
Figure 8: LEH and BSC network. Left Mid March 2008 Right Mid September 2008
@@
Table 2 provides a concise summary of the logic above. When both node- and system-level LCMs are on the rise, more instability is introduced coming from individual institutions. Conversely, a declining trend at the node-level while the system-level LCM is increasing demands a closer look at the specific firm. Such divergence implies the firm is treated independently from the network and can imply different conditions: if the firm is struggling, portfolio detachment is triggered; if the firm is in a good shape, it is considered a safe investment opportunity. Similarly, when the system-level LCM is decreasing while the node-level is increasing, it holds implications based on the firm’s condition; a struggling firm means a high likelihood of instability, while a healthy one becomes densely interconnected within the network, drawing heavy reliance from investors. This interconnectedness elevates the risk of contagion under a shock scenario to these robust entities.
Table 2: Node- and system-level LCM dynamics
@@ 5 Conclusion
This paper presents a framework for constructing a weighted directed network, computing the absolute centrality measure, and statistically validating it. This offers a robust approach to assess contagion risk and vulnerability within financial systems. The Leontief centrality measure serves as a reliable proxy for contagion risk, allowing for meaningful comparisons across different time periods. To further enhance analysis of the network, we derive the distribution of the centrality measure, enabling the computation of confidence intervals and hypothesis testing. We show the alignment of the theoretical and empirical centrality distributions with the true centrality measure using simulated data. Leveraging real world financial data, CDX spread and equity tick data, we construct financial networks and evaluate the centrality of the system. During distressed years, we observe higher LCM values, underscoring the measure’s ability to capture systemic risk within the network and assess the investors’ beliefs. LCM can be used to form clusters of firms based on their susceptibility to contagion risk, thereby being useful for factor investing strategies and risk management practices. Central institutions through this measure can be identified to determine which firms require more rigorous oversight.
Future research can expand the scope of this work to, e.g., incorporate different numbers of lags in building the network. Such a construction would introduce a multilayered centrality metric. This extension could offer a more comprehensive understanding of the complex dynamics and interdependencies observed in financial data.
@@ Appendix A Proofs
A.1 Proof of Pair-Level LCM Distribution (Theorem 3)
@@
Under the assumptions of this theorem, the noise of the model $\Delta$ in Equation 4 follows a multivariate normal distribution. In fact, due to the linear relation between the (first-order approximation of the) pair-level LCM $(I-\hat{A})^{-1}_{ij}$ and the noise $\Delta$, it trivially follows that the pair-level LCM must also follow a normal distribution. Thus, to complete this proof we only need to consider the mean and variance of this linear form.
@@
Second, consider the variance of the noise, i.e., $var(\Delta_{\cdot j})=\sigma^{2}_{j}(X^{\top}X)^{-1}$ where $\sigma_{j}$ is the standard deviation of the noise of the $j^{th}$ time series by standard OLS arguments. Therefore, by the first-order Taylor expansion specified above,
@@
$=\left[(I-A)^{-1}_{\cdot j}\right]^{\top}\left[(I-A)^{-1}_{i\cdot}\Sigma_{lk}^{T}\left[(I-A)^{-1}_{i\cdot}\right]^{\top}\right]_{lk}(I-A)^{-1}_{\cdot j}$
where $\Sigma_{lk}^{T}=\rho_{lk}\sigma_{l}\sigma_{k}Q_{T}^{-1}$ provides the covariance between $\Delta_{\cdot l}$ and $\Delta_{\cdot k}$ (with multiplier $\sqrt{T}$). For this formulation we specify $Q_{T}=T^{-1}X^{\top}X$. As $Q=\lim_{T\to\infty}Q_{T}$, the result follows.
A.2 Proof of Node-Level LCM Distribution (Corollary 4)
@@
A.3 Proof of Test Statistic’s Distribution (Corollary 5)
@@
Appendix B Additional Details and Figures for Section 4.1
@@
Figure 10: Validated, not validated and shaded confidence interval of CDX LCM.
Appendix C Details of the Data Set for Section 4.2

### v2

@@ 4.2 Equity Tick Market Contagion
The choice of data frequency is a critical consideration in network analysis. To ensure that dependencies are captured in the model, we must strike the right balance between a frequency that is not too low or too high. Given that stock prices swiftly adjust to market changes, especially to similar stock price fluctuations, and trading occurs at high frequency to capitalize on these changes, we opt for tick data. We use tick log return data for 20 Financials sector companies (provided in Table 3 of Appendix D) as well as both Lehman Brothers (LEH) and Bear Stearns (BSC). The tick data is obtained from Refinitiv, collected at one-second intervals, and spans from January 1, 2005 to December 31, 2010 aligning with Martinez-Jaramillo et al., (2014).
@@
In Figure 8, we focus on the period surrounding the 2008 Financial Crisis to gain a deeper understanding of the centrality measure’s behavior in the years preceding, during, and after the turmoil. We consider two versions of the system-level KB: (i) for the network including LEH and BSC, for which we run the analysis for as long as the data is available (October 8, 2008 indicated by a red dashed line in the figure) and (ii) for the network excluding these two firms. In both versions, we observe an upward trend starting from early 2007 with the peak KB reached in August 2008. A gradual decline is then observed when excluding LEH and BSC. This signifies that the contagion risk increased until reaching the peak of turmoil, subsequently decreasing after government interventions and as the economy and markets slowly returned to normalcy. This system-level trend closely aligns with the timeline of the financial crisis: the acquisition of Bear Sterns by JPMorgan on March 16, 2008, and the subsequent default of Lehman Brothers on September 15, 2008. The Wall Street Bailout law was passed on October 3, 2008, which included the \$700 billion Troubled Asset Relief Program (TARP) aimed at rescuing banks from defaulting.
@@
We plot the node-level KB for three specific firms (American International Group (AIG), JPMorgan (JPM) and Goldman Sachs (GS)) alongside Lehman Brothers and Bear Stearns in Figure 8. We note that the node-level KB for LEH and BSC have similar behavior to the system-level KB until May 2008. Afterwards, they start to decline, due to which we experience a slight decline in the system-level KB as well. The node-level drop is aligned with recognition of the substantial risk these two companies carried by investors. Due to this market behavior, investors treated these companies idiosyncratically thus reducing the spillover effects to the rest of the system from shocks they might experience. This is why, despite the systemic contagion risk rising, the KB for LEH and BSC move inversely. This detachment is also reflected in the network structure. Figure 9 shows the interdependencies between LEH and BSC and all other firms for mid March and September 2008 time windows. As is evident, the connectivity of these two companies diminishes in both quantity and significance during their corresponding drop in their KB.
@@
Analyzing node-level KBs reveals some similarities and differences between the three tickers, AIG, JPM and GS, and the delisted ones. We notice that AIG, prominent in the insurance sector, exhibits a similar behavior to LEH and BSC; its KB took a downward turn following the bankruptcy and merger of LEH and BSC respectively. During the crisis, \$182 billion was injected into AIG to prevent its default by the Treasury and Federal Reserve Bank of New York, as it was believed that its collapse would jeopardize the integrity of other firms. Essentially, AIG came under government ownership and was safeguarded, which is reflected in its KB as it reaches nearly zero by early 2009. On the other hand, JPM and GS, for which the KB consistently increases even after October 8, 2008, diverge from the system-level KB. As these two companies were considered safer by investors, they maintained investments despite the systemic turmoil. However, under a shock scenarios to these companies, there would have been a high contagion risk. Looking into the stock prices of these tickers, we observe a significant decline in AIG’s price, whereas JPM and GS experience a decrease but recover shortly afterwards. These patterns are aligned with the fluctuations seen in the KB, indicating its ability to mirror the dynamics of the market.
Figure 8: Node- and system-level KB
@@
Figure 9: Incoming and outgoing edges of LEH and BSC network. Left Mid March 2008 Right Mid September 2008. The blue colored nodes are LEH and BSC
@@
Table 2 provides a concise summary of the logic above. When both node- and system-level KBs are on the rise, more instability is introduced coming from individual institutions. Conversely, a declining trend at the node-level while the system-level KB is increasing demands a closer look at the specific firm. Such divergence implies the firm is treated independently from the network and can imply different conditions: if the firm is struggling, portfolio detachment is triggered; if the firm is in a good shape, it is considered a safe investment opportunity. Similarly, when the system-level KB is decreasing while the node-level is increasing, it holds implications based on the firm’s condition; a struggling firm means a high likelihood of instability, while a healthy one becomes densely interconnected within the network, drawing heavy reliance from investors. This interconnectedness elevates the risk of contagion under a shock scenario to these robust entities.
Table 2: Node- and system-level KB dynamics
@@ 5 Conclusion
This paper presents a framework for constructing a weighted directed network, computing the centrality measure and statistically validating it. This offers a robust approach to assess contagion risk and vulnerability within financial systems. The KB centrality measure serves as a reliable proxy for contagion risk, allowing for meaningful comparisons across different time periods. To further enhance analysis of the network, we derive the distribution of the centrality measure, enabling the computation of confidence intervals and hypothesis testing. We show the alignment of the theoretical and empirical centrality distributions with the true centrality measure using simulated data. Leveraging real world financial data, CDX spread and equity tick data, we construct financial networks and evaluate the centrality of the system. During distressed years, we observe higher KB values, underscoring the measure’s ability to capture systemic risk within the network and assess the investors’ beliefs. KB can be used to form clusters of firms based on their susceptibility to contagion risk, thereby being useful for factor investing strategies (Packham, , 2024) and risk management practices. Central institutions through this measure can be identified to determine which firms require more rigorous oversight.
Future research can broaden the scope of this work by exploring alternative network construction approaches, experimenting with different numbers of lags or markets (Abinash Mishra and Chakrabarti, , 2022), and incorporating these variations into the network’s structure. Such advancements could lead to the development of a multilayered centrality metric. From a financial perspective, applying this enhanced centrality measure to calculate risk metrics, such as Value at Risk, Expected Shortfall, CoVaR (Adrian and Brunnermeier, , 2016), and MES (Acharya et al., , 2016), would be particularly valuable. This extension could offer a more comprehensive understanding of the complex dynamics and interdependencies observed in financial data.
Disclosure of interest
The authors have no conflicting interests to declare.
Declaration of funding
No funding was received for this work.
@@ Appendix A Proofs
A.1 Proof of Pair-Level KB Distribution (Theorem 3)
@@
Under the assumptions of this theorem, the noise of the model $\Delta$ in Equation B follows a multivariate normal distribution. In fact, due to the linear relation between the (first-order approximation of the) pair-level KB $(I-\hat{A})^{-1}_{ij}$ and the noise $\Delta$, it trivially follows that the pair-level KB must also follow a normal distribution. Thus, to complete this proof we only need to consider the mean and variance of this linear form.
@@
Second, consider the variance of the noise. By the first-order Taylor expansion specified above,
@@
$=\left[(I-A)^{-1}_{\cdot j}\right]^{\top}\left[(I-A)^{-1}_{i\cdot}\Sigma_{lk}^{T}\left[(I-A)^{-1}_{i\cdot}\right]^{\top}\right]_{lk}(I-A)^{-1}_{\cdot j}\,,$
where $\Sigma_{lk}^{T}$ provides the covariance between $\Delta_{\cdot l}$ and $\Delta_{\cdot k}$ (with multiplier $\sqrt{T}$).
A.2 Proof of Node-Level KB Distribution (Corollary 4)
@@
A.3 Proof of Test Statistic’s Distribution (Corollary 5, Corollary 6)
@@
Appendix B Vector Autoregressive System (Section 4)
As we calibrate the adjacency matrix as the OLS estimates, we recall this construction for a generic multivariate model. Given two matrices $X,Y\in\mathbb{R}^{T\times M}$ (over real numbers), where $T$ represents the number of time steps and $M$ denotes the number of variables, and an iid noise term $\mathcal{E}\in\mathbb{R}^{T\times M}$, we can adopt a linear regression framework as below:
$\begin{gathered}Y=XA+\mathcal{E}\,,\\ \hat{A}=(X^{\top}X)^{-1}X^{\top}Y\,,\end{gathered}$
where $A$ is the true coefficient matrix that governs the relationships between the elements of $X$ and $Y$. On the other hand, $\hat{A}$ represents the estimated coefficient matrix, which is obtained from the regression analysis. Define $Q:=\lim_{T\to\infty}T^{-1}X^{\top}X$ (with limit taken in probability as the number of data points grows to infinity) and set $\Sigma_{lk}:=\rho_{lk}\sigma_{l}\sigma_{k}Q^{-1}$ for each $l,k$ where $\rho_{lk}$ denotes the correlation coefficient between time series $l,k$ and $\sigma_{k}$ denotes the standard deviation of time series $k$ residuals.
Appendix C Additional Details and Figures for Section 4.1
@@ Appendix B Vector Autoregressive System (Section 4)
Figure 11: Validated, not validated and shaded confidence interval of CDX KB.
Appendix D Details of the Data Set for Section 4.2